OpenStudy (anonymous):

classify the triangle based on side lenghts 5,15,19. A.right B.acute c.obtuse d. no triangle can be formed with given side lengths

OpenStudy (campbell_st):

a triangle can be formed and you can use pythagoras' theorem if c^2 < a^2 + b^2 the triangle is acute angled c^2 = a^2 + b^2 the triangle is right angled c^2 > a^2 + b^2 the triangle is obtuse angled then 19^2 = 15^2 + 5^2 361 > 250 the triangle is obtuse angled.